OpenStudy (anonymous):

Which expression is the first step you take when finding the quotient of seven over eight division symbol seven over sixteen? eight-sevenths times seven-sixteenths seven-eighths times seven-sixteenths seven-eighths times sixteen-sevenths eight-sevenths times sixteen-sevenths

OpenStudy (agent0smith):

\[\Large \frac{ 7 }{ 8 } \div \frac{ 7 }{ 16 }\] dividing by a fraction is the same as multiplying by the reciprocal \[\Large \frac{ 7 }{ 8 } \times \frac{ 16 }{ 7 }\]